Your talk title

Your name Date of talk

Outline

1 This is a section for your talk 2 Here’s another section for your talk

1 This is a section for your talk

• You go on to make many fascinating points.

• And more fascinating points.

Notes for Slide 2

Here’s some discussion of the fascinating points you made.

2 Here’s another section for your talk

• Whoa!! You are making more points!!

• And even more, using overlays!!

2 Here’s another section for your talk

• Whoa!! You are making more points!!

• And even more, using overlays!!

Notes for Slide 3

Now these extra points you made certainly require some discussion.